OpenStudy (anonymous):

Please help!! Will give medal! Generate the first 5 terms of this sequence F(1)=2 and f(2)=3, f(n) =f (n-1)+f(n-2), for n>2 A. 2,3,5,7,9 B. 2,3,4,7,11 C. 2,3,5,9,11 D. 2,3,5,8,13

OpenStudy (anonymous):

Well, to find the 3rd term, n = 3, you'd want f(3-1) + f(3-2) and that would be your 3rd term in the sequence. Continuing from there, the 4th would be f(4-1) + f(4-2). In essence, check what the sequence is asking for :). f(2) + f(1) = f(3) You are given f(2) and f(1). So what is f(2) + f(1) ?

OpenStudy (princeharryyy):

f(3)=f(1)+f(2)=2+3 =5; f(4) = f2+f3=3+5 = 8; f(5) = f4+f3 =8+5 = 13; ans d) @WhiteFlower :D

OpenStudy (anonymous):

+1 for right answer, -1 for not stepping her through the concept :P @princeharryyy

OpenStudy (anonymous):

Do you think you can help me? @princeharryyy

OpenStudy (princeharryyy):

no big deal in concept , every next term is the sum of last two terms @adam.bloedel hehehe :p

OpenStudy (princeharryyy):

maybe @Ccarey24

OpenStudy (anonymous):

That's easy for us to understand, but she might not know what it meant. Best bet is to always explain how and why. Excellent job showing steps though. :)

OpenStudy (princeharryyy):

tnks @adam.bloedel

OpenStudy (anonymous):

Thank you @princeharryyy for showing me the steps and letting me know where I went wrong! And also, thank you @adam.bloedel for also helping!

OpenStudy (princeharryyy):

u r welcome @WhiteFlower